In Desmos: type in the equations below and match the graph with each equation


y= square root of x + 2 -1

y= square root of x -3+1

y= 3 square root of x-3+2

y= 3 square root of x+ 2 + 3


1
Expert's answer
2021-03-16T07:15:25-0400

■ y = "\sqrt{x+2-1}" = "\sqrt{x+1}"



■ y = "\sqrt{x-3+1}" = "\sqrt{x-2}"



■ y = 3 "\sqrt{x-3+2}" = 3 "\sqrt{x-1}"


■ y = 3 "\sqrt{x+2+3}" = 3"\sqrt{x+5}"